OpenStudy (anonymous):

The sides of an equilateral triangle are increasing at the rate of 2 ft/min. Find the rate of change of the Perimeter and Area of the triangle when the side is 3 ft. please help me with this one....

OpenStudy (anonymous):

optimization?

OpenStudy (anonymous):

i dont know.. :(

OpenStudy (anonymous):

what class are you in?

OpenStudy (anonymous):

it's a lesson in my calculus subject...

OpenStudy (anonymous):

yeah it is optimization, I would recommend watching videos from patrickjmt on youtube. http://patrickjmt.com/optimization-problem-1/

zepdrix (zepdrix):



zepdrix (zepdrix):

So they gave us this information. Let's think about the perimeter as a function of the side length.\[\Large\rm P(x)=3x\]Right?

zepdrix (zepdrix):

I added up the 3 sides, x+x+x, that step ok? :o

OpenStudy (anonymous):

then...?

zepdrix (zepdrix):

Ummm we want to take the derivative of our function `with respect to time`. So what will happen with our x when we do that?

zepdrix (zepdrix):

Remember, we're not differentiating with respect to x.

zepdrix (zepdrix):

\[\Large\rm P'(x)=?\]

zepdrix (zepdrix):

Where you at jesssss >.< You can do itttttt

zepdrix (zepdrix):

If this is confusing, we can switch to Leibniz notation.\[\Large\rm P=3x\]And we want to find:\[\Large\rm \frac{dP}{dt}\]

OpenStudy (anonymous):

3 dx?

zepdrix (zepdrix):

Ok good!\[\Large\rm P'(x)=3(1)x'\]

OpenStudy (anonymous):

P' <--- is this the rate of change of the perimeter?

zepdrix (zepdrix):

Yes, \[\Large\rm P'=\frac{dP}{dt}\]the rate of change of perimeter over time

OpenStudy (anonymous):

so i'll just substitute 2 in P'(x) = 3(1)x' ...?

zepdrix (zepdrix):

x'=2 dx/dt=2 This is how the length is changing over time. So plug that in first of all :O

zepdrix (zepdrix):

And then figure out P'(3). The rate of change of perimeter over time when the side length is 3.

zepdrix (zepdrix):

I hope I don't trick you up the way I worded that >.< since there are no x's to plug in for...

OpenStudy (anonymous):

P'(3) is 6...?

zepdrix (zepdrix):

Ok great! :) So we've found the rate of change of the perimeter. It turns out it's independent of the side lengths. it only depends on the rate of change of the side.

zepdrix (zepdrix):

So for area, ummmmm area of an equilateral triangle... i better look that one up >.<

zepdrix (zepdrix):

Google is telling me that it's,\[\Large\rm A(x)=\frac{\sqrt3}{4}x^2\](We don't want another variable, so we needed a formula that doesn't involve height).

zepdrix (zepdrix):

I suppose we could have used Heron's Formula to derive that, but Google saves time >.< heh

zepdrix (zepdrix):

So what do you think we'll do here? c:

OpenStudy (anonymous):

differentiate the Area, then plug in x' ... then solve for A'(3)...? :)

zepdrix (zepdrix):

Yah good c: Since we have a square on x, it looks like our area function will end up depending on both x and x'. So we'll have more stuff to plug in this time.

zepdrix (zepdrix):

Do eet! >.< tell me what you get!

OpenStudy (anonymous):

\[3\sqrt{3}\] ...?

zepdrix (zepdrix):

Mmm good that's what I'm coming up with also \c:/ yay team! Hopefully we didn't make any boo boo's along the way :U Seems pretty straight forward.

OpenStudy (anonymous):

what is the unit for A'...? ft^2/min. ..?

OpenStudy (anonymous):

ft/min for P' ...?

zepdrix (zepdrix):

Ummm yah I guess that would make sense. I'm trying to remember how the units pop out when we do it the legit way... Oh I guess plugging in the values is where we get the dimensions from.\[\Large\rm A'(3)=\frac{\sqrt3}{2}\left(3ft\right)\left(2\frac{ft}{\sec}\right)\]

zepdrix (zepdrix):

/min* soz

OpenStudy (anonymous):

thanks for your help! :)

zepdrix (zepdrix):

\c:/

OpenStudy (isaiah.feynman):

I missed the show! :(